4.- (25) Find it) two ways: a) in the time domain the Laplace Transform 6) using t=0 sec + 10 cost Volts 12 ww lovolts IH
Answer #1

The circuit given lo cost 14 10V to find Apply super position the dem response Consider 10V de source ilts 10=Riti dilat TH 1i trannient K ett zket i(Hz No Cos (6-45°) + Ké o) + két? to find K value consider i(t) at tao i Col=0 COM45° + Ke к (t = c6 Apply Laplace trantam 10 coat effor wat] = 1 tout ict) + di locost +10 L[k] = kls dt IOS ICS) [1+5] S 1+s? IOS + ICS) S(+1)ilt) = 10-15ét + 5(cost+ Bint) iltl - | 0 – 15 - + Tocos (t-45°) (cost+ sint) 1 = 2 Te I cost.c0348°+ Bfnt.sings] Te [coj[t-4